OpenStudy (anonymous):

What are all the kinds of indeterminate forms you can have with infinity and 0?

OpenStudy (anonymous):

i dont remeber all, 0/0;inf/inf;1^inf;inf-inf;0^inf

OpenStudy (anonymous):

Well, there are an infinite number of indeterminate forms...those are definitely some of the most common noes though. However, I don't believe \(1^\infty\) is indeterminate.

OpenStudy (anonymous):

@yakeyglee it's indeterminate for sure!

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=1%5Einf

OpenStudy (anonymous):

Yeah, you're right. Whoops! For instance, on one hand, we have \[\lim_{x \rightarrow \infty} \left( 1+ \frac{1}{n}\right)^n=e,\]and on the other, \[\lim_{x \rightarrow \infty} \left( 1+ \frac{1}{n^2}\right)^n=1.\]

OpenStudy (anonymous):